OpenStudy (anonymous):

Algebra Help!!! determine the discriminant of the equation. How many real solutions does it have? x^2-5x+7=0?

OpenStudy (anonymous):

always use the equation b^2 - 4ac x²-5x+7=0 therefore a=1 b=-5 c=7 therefore the answer would be (-5)²-4(1)(7) =25-28 =-3 the answer is -3

OpenStudy (anonymous):

Hope this helps :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

medals please

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

would you mind helping me on another one

OpenStudy (anonymous):

what are the solutions?\[\frac{ 1 }{ 2 }x^{2}+2x+3=0\]

OpenStudy (anonymous):

ok hang on a second

OpenStudy (anonymous):

1/2x²+2x+3=0 2(1/2x²+2x+3)=0*2 x²+4x+6=0 x²+4x=-6 【(4/2)²=4】 x²+4x+4=-6+4 【completing the square】 (x+2)*(x+2) =-2 (x+2)² =-2 x=(+√-2)-2 x=(-√-2)-2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

its okay i hope you understand it that is important to you :)